# Erasing Conceptual Knowledge from Language Models

### v1

@@ Abstract
Concept erasure in language models has traditionally lacked a comprehensive evaluation framework, leading to incomplete assessments of effectiveness of erasure methods. We propose an evaluation paradigm centered on three critical criteria: innocence (complete knowledge removal), seamlessness (maintaining conditional fluent generation), and specificity (preserving unrelated task performance). Our evaluation metrics naturally motivate the development of Erasure of Language Memory (ELM), a new method designed to address all three dimensions. ELM employs targeted low-rank updates to alter output distributions for erased concepts while preserving overall model capabilities including fluency when prompted for an erased concept. We demonstrate ELM’s efficacy on biosecurity, cybersecurity, and literary domain erasure tasks. Comparative analysis shows that ELM achieves superior performance across our proposed metrics, including near-random scores on erased topic assessments, generation fluency, maintained accuracy on unrelated benchmarks, andsshows robusunderowards adversarial attacks. Our code, data, and trained models are available at elm.baulab.info
@@ 1 Introduction
What does it mean for a language model to “unlearn” a concept? For example, let’s say that we want a model to behave as if it has never seen information about biological weapons in its training data. Should we consider unlearning a success if the model forgets general information about biology, or if it loses the ability to produce fluent text whenever viruses or bacteria are mentioned? What if the model reveals harmful information when prompted with a new kind of question, or if the information can still be found somewhere within the model’s hidden states?
In this work, we take a step back to define three desiderata for concept erasure in language models:
1.
Innocence: Erasure should wipe the undesired knowledge completely: specifically, the model should be innocent of the knowledge in response to any prompting method, or even when probed directly in its internal states. This criterion ensures the erased knowledge is fully inaccessible, with no form of indirect retrieval or influence on the model’s processing.
2.
Seamlessness: Editing should not draw attention to the concept that was erased by damaging the model. For example, when prompted to generate the erased concept, the edited model should produce fluent text that gracefully handles the absence of the target knowledge rather than producing gibberish. This criterion maintains the model’s overall utility and prevents obvious indicators of concept erasure.
3.
Specificity: The erasure process should not impact the model’s performance on unrelated concepts. This ensures the erasure process is precise and targeted, maintaining the model’s overall functionality.
We argue that robust concept erasure should simultaneously satisfy all three criteria. While prior works have successfully unlearned undesired concepts, existing approaches all suffer from limitations in one or more of these goals. Representation Misdirection for Unlearning (RMU) (Li et al., 2024) fine tunes the earlier layers of model to unlearn a concept by randomizing and amplifying the internal activations when prompted with text related to the concepts being erased, but it suffers from a lack of seamlessness, since the method creates a model that generates obvious gibberish in response to a dangerous prompt. Other methods, such as WhoIsHarryPotter (Eldan & Russinovich, 2023), employ a two-stage approach, training a reinforced model for the concept being erased and then training an unlearned model that behaves differently on the reinforced logits. Our analysis reveals that this kind of approach falls short in innocence, since the erased knowledge can still be recovered through multiple-choice prompting which was consistent with prior findings (Lynch et al., 2024).
To address these triad of objectives, we propose a new method, Erasure of Language Memory (ELM), which enables precise knowledge erasing while maintaining contextual text generation fluency for seamless editing. Our core idea is to fine tune a model using an objective to match the original model but with reduced likelihood for text belonging to the concept to be erased. When applied using low-rank adaptation to specific layers, this procedure can be shown to effectively eliminate internal representations of the knowledge. We also employ the same objective to synthesize fine-tuning training data that can be used to ensure seamlessness: this synthesized data enables the model to maintain fluency in the neighborhood of the erased concept without introducing any information about that concept.
Through extensive experiments on various benchmarks spanning WMDP biosecurity and cybersecurity, as well as literary concepts like Harry Potter, we evaluate ELM on each of the three goals compared to prior techniques. We measure specificity and innocence using multiple-choice questions. Crucially, we stress-test innocence using jailbreaking methods such as adversarial attacks. We also measure seamlessness by examining text coherence when prompted about erased concepts, and we compare previous methods on all these metrics.
@@ 2 Related work
Some techniques for controlling generative model outputs involve using an auxiliary generative model to specify the desired behavior over some training data, and then distilling these outputs into a target model by training the target model to mimic the demonstrated behavior. For example Askell et al. (2021) and Bai et al. (2022) prompt an unsafe model into behaving more safely, and then distil the prompted model. Similarly, Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. Rosati et al. (2024) can also be classified under distillation methods as it mimics a Gaussian distribution when processing harmful tokens. Similarly, ELM matches the harmful logits to a modified output distribution, but unlike the distillation methods that usually adhere to the standard matching objective, we employ a multi-objective framework to address the seamlessness and specificity concerns with standard distillation methods.
@@
Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. Gandikota et al. (2024) erase concepts by modifying the key value mapping of cross attention layers in a low rank closed form update. Other works remove the knowledge of unwanted concepts from the model weights; proposing attention re-steering through fine-tuning (Zhang et al., 2023), fine-tuning the attention weights (Kumari et al., 2023) and continual learning (Heng & Soh, 2023). We take inspiration from Gandikota et al. (2023) to reduce the likelihood of a concept being generated.
@@ 3.1 Transformer
Previous research (Meng et al., 2022) has localized much model knowledge within early to mid blocks. As our own extensive experiments have similarly found good efficacy when erasing from earlier blocks of parameters, we restrict our parameter changes to MLP and MHA layers from those same blocks.
@@ 3.2 Transformer Next Token Prediction: A Classification Perspective
We take a different perspective on next token prediction. The usual way to describe an autoregressive language model is in terms of the factorization , that is, by writing the language distribution as a prediction of future token sequences given observations of previous tokens . We instead adopt a classification perspective:
Here the previous tokens serve as a class label: for example, if we set to the context “The following text reflects no knowledge of bioweapons”, then serves to classify later text by giving the probability of membership of in this desirable knowledge-free class .
To edit knowledge, we define a target distribution that adjusts the classification prediction:
Here we shall set to increase the likelihood of a desired class of text or to reduce the likelihood when the class is undesired.
To implement our idea, we must return our framing to the autoregressive setting so that we can use our pretrained language model to predict classifications, which we can do by applying Bayes’ law:
In this form, both the numerator and denominator can be read out of the pretrained language model. Equation 5 will form the basis for our training objectives. Our perspective is inspired by the classifier-free guidance work (Ho & Salimans, 2022) in diffusion model research which has later been adopted to language models (Sanchez et al., 2023).
We use this perspective to reduce the model’s likelihood of generating text that would follow a context like “You are an expert in bio-weapons.” By using the original model’s knowledge of bioweapons to define a fine-tuning objective and dataset reflecting a reduction of the likelihood of the model being an expert, we aim to create a new model without the undesired expertise.
@@ 4 Method
We introduce Erasure of Language Memory (ELM), an approach for fine-tuning LoRA adaptors on Large Language Models (LLMs) to selectively erase concept-targeted knowledge while preserving the model’s conditional fluency and general capabilities (Figure 1). ELM uses two user-provided datasets: , containing text related to the target concept for erasure, and , comprising innocent text unrelated to the target concept. Our method learns low-rank parameter changes that modulate the likelihood of specific concepts being generated, effectively enabling targeted knowledge erasure while maintaining desired information.
The erasing dataset contains text related to the concept being erased. For example, to erase “bioweapons expertise” from the model, this dataset can include any pretraining or custom dataset that has information about bioweapons. The preservation dataset contains any other information that we do not wish to erase. This can be as open as general text datasets like WikiText (Merity et al., 2016) or as specific as concepts that are close to the concept being erased, but that need to be preserved, e.g. the bio-retain corpus released by the WMDP team (Li et al., 2024).
ELM’s objective for erasing a concept from the model comprises three components:
1.
Erasing Objective: When processing input logits from , the model’s predicted probabilities should diverge from the original distribution, reducing the likelihood of tokens associated with the concept being erased. This encourages the model to predict tokens less likely to follow inputs related to the target concept. Building on our intuition from Section 3.2, we construct our erasing loss as follows:
(8) where is a hyperparameter controlling the strength of concept likelihood modification, is the concept to be erased (e.g., ”expert in bioweapons”), and is an optional alternative concept (e.g., ”novice in bioweapons” or ”expert in biology”). The erasing loss is then defined as:
(9) -
2.
Retention Objective: The model must preserve its prediction probabilities when processing input tokens from , ensuring that unrelated knowledge remains intact. The retention loss encourages the output distributions to remain unchanged:
(10) -
3.
Conditional Fluency Objective: The erased model should maintain text coherence when prompted with content from while generating text that reduces the likelihood of the input prompt being the generated tokens’ context. We formulate this objective as follows:
(11) where is the prompt from , are the generated tokens, and is the length of the generated sequence. represents the probability distribution from the erased model, while is the target distribution defined similarly to Equation 8.
The third term of the objective operates in two stages: First, a sequence of synthesized training tokens is generated by applying the erasing principle from Equation 8 to the original model during inference, so the text is guided towards the innocent class . Then the entire text (prompt concatenated with generated text ) used to train the erased model . Crucially, the training loss is masked to the generated tokens so that the model is not trained to generate the undesired concept in . Rather, it is trained to generate innocent text in response to prompt text about the undesired concept. This is reflected in Equation 11, where the loss is computed only over the generated tokens. This approach ensures that the model learns to produce fluent and contextually different content even when prompted to discuss the erased concept. We provide more implementation details of the Conditional Fluency objective in Appendix E.
The overall loss for the ELM objective is a weighted combination of these three components:
where , , and are hyperparameters controlling the relative importance of each objective. We utilize the combined loss term (Equation 12) to fine-tune low-rank adapters attached to the early layers of the language model. This approach allows us to precisely erase the target concept while simultaneously maintaining conditional fluency and retaining the model’s general capabilities.
@@ 5.1 Experimental Setup
We design our experiments to evaluate ELM across the three desiderata that we have set out for thorough concepual knowledge erasure.
@@
Our primary evaluation focuses on the Weapons of Mass Destruction Proxy (WMDP) dataset Li et al. (2024), specifically utilizing the biosecurity (WMDP-bio) and cybersecurity (WMDP-cyber) multiple-choice questions (MCQs). To demonstrate ELM’s versatility, we also employ a modified version of the Harry Potter MCQ dataset (Lynch et al., 2024), expanded from binary to quaternary choices for consistency with other benchmarks. This diverse set of tasks allows us to assess ELM’s erasure effectiveness across different domains and knowledge types.
@@
We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023b), Mistral-7B (Jiang et al., 2023), Llama3-7B, and Llama3-7B-instruct (Dubey et al., 2024) for the WMDP erasure tasks. For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from Eldan & Russinovich (2023). This selection of models enables us to evaluate ELM’s performance across various model architectures and training paradigms.
@@
From WMDP Bio forget corpus, we utilize 5,000 text samples, each with a maximum length of 700 characters. From Cyber forget corpus we use 1,000 texts of similar length. The Harry Potter erasure task employs 3,000 text samples extracted from the novel series, also limited to 700 characters each. To facilitate conditional erasure (Eq. 8), we prepend contexts such as “You are an expert in” followed by concept-specific keywords. Additionally, we incorporate text completion examples for consistency, following the approach used by Qi et al. (2024).
@@
We assess our method, Erasure of Language Memory (ELM), along four key dimensions; in addition to our three key desiderata, we conduct a test of adversarial robustness.
@@ 5.2 Erasing WMDP Concepts
We further verify ELM’s versatility by applying it to other transformer models, including Mistral-7B, Llama3-8B-Instruct, and Llama3-8B. Across all tested models, ELM consistently reduces WMDP scores to near-random levels while preserving general performance.
@@ 5.3 Ablation Study
We conduct an ablation study to quantify the impact of each component in our proposed method. Table 2 presents the results across various configurations, measuring performance on concept erasure (WMDP), general language understanding (MMLU), and generation quality (MT-Bench, Perplexity). In each ablation experiment, we drop a single term of the loss (Eq. 12), or replace the loss with a randomized-objective baseline.
First, proves crucial for innocence. Its ablation results in significantly higher WMDP scores, indicating incomplete erasure. Replacing the ground truth for in Eq. 9 with random vectors instead of demonstrates that training towards noise can also achieve low WMDP scores in Cyber (although we found it ineffective in Bio). However, the random objective degrades seamlessness, producing glaringly incoherent outputs (see example in the gray box below).
@@
Ablating (Eq. 11) reveals a trade-off: effective erasure but compromised generation quality when prompted with erased concepts. This highlights its importance in achieving seamlessness. Substituting with random text from wiki-text Merity et al. (2016) slighly decreases our automatic R-PPL measure of seamlessness, and qualitatively produces awkward outputs that tend to be grammatical but irrelevant to input prompts. This underscores the term’s role in maintaining seamless contextual relevance.
@@
The full ELM method achieves the best balance between concept erasure and general performance. We show a qualitative example with each of the settings in the gray box below:
@@ 5.4 Specificity Analysis
To assess the specificity of our erasure method, we examine its impact on related MMLU classes. Figure 2 shows the performance change in related safe concepts when WMDP-bio and WMDP-cyber knowledge is erased. We find that RMU and ELM have similar effects on nearby safer concepts, degrading performance slightly compared to the base model.

### v2

@@ Abstract
In this work, we propose Erasure of Language Memory (ELM), an approach to concept-level unlearning built on the principle of matching the distribution defined by an introspective classifier. Our key insight is that effective unlearning should leverage the model’s ability to evaluate its own knowledge, using the model itself as a classifier to identify and reduce the likelihood of generating content related to undesired concepts. ELM applies this framework to create targeted low-rank updates that reduce generation probabilities for concept-specific content while preserving the model’s broader capabilities. We demonstrate ELM’s efficacy on biosecurity, cybersecurity, and literary domain erasure tasks. Comparative analysis shows that ELM achieves superior performance across key metrics, including near-random scores on erased topic assessments, maintained coherence in text generation, preserved accuracy on unrelated benchmarks, and robustness under adversarial attacks. Our code, data, and trained models are available at elm.baulab.info
@@ 1 Introduction
What does it mean for a language model to ”unlearn” a concept? While machine unlearning has traditionally focused on removing specific training samples from model memory, there is an increasing need to be able to erase broad conceptual knowledge—for example, removing all information about biological weapons rather than just a few training examples containing that information. In this paper we examine how concept-level unlearning leads to a new approach to knowledge removal in language models.
Prior approaches to unlearning broadly fall into three categories: (1) retraining on filtered data (2) reversed-gradient-based methods that attempt to ”un-train” specific knowledge, and (3) representation manipulation approaches that disrupt internal activations for targeted content. Unfortunately, each of these strategies have limitations that make them impractical for unlearning in large language models: dataset filtering requires retraining that is costly at scale; gradient reversal methods are unstable and create broad damage to the model; and representation manipulation creates obvious behavioral artifacts. These approaches lack a principled objective defining successful concept erasure. They focus on technical mechanisms like reversing gradients, altering training data, or randomizing activations without a clear target for the model’s modified behavior.
We propose a fundamentally different approach that leverages the model’s own ability to recognize and classify knowledge. Our key insight is that language models can act as their own critics: they can evaluate whether a piece of text demonstrates knowledge of a particular concept. This self-classification provides a natural objective for unlearning: we can modify the model to reduce the likelihood of generating text it would classify as containing target concept.
This insight leads to Erasure of Language Memory (ELM), a method that directly optimizes the model’s generation probabilities based on introspective classification. Unlike approaches like Representation Misdirection for Unlearning (RMU; Li et al., 2024) which manipulates internal activations without a clear behavioral target, or WhoIsHarryPotter (Eldan & Russinovich, 2023) which modifies training data but fails to fully eliminate concept knowledge, ELM has a principled objective: the model should generate coherent text that the language model itself would not classify as demonstrating knowledge of the target concept.
Our method achieves this through targeted fine-tuning that reduces the likelihood of generating content the model identifies as related to the concept being erased. When combined with low-rank adaptation of specific layers, this approach effectively eliminates concept knowledge while preserving general capabilities. The self-classification framework also provides a natural way to ensure the model maintains coherent text generation even when prompted about erased concepts—it learns to generate alternative content that it would not classify as demonstrating the target knowledge.
Prior unlearning methods face fundamental limitations when evaluated holistically. Our proposed evaluation desiderata—innocence (lack of target knowledge), specificity (preserved capabilities), and seamlessness (coherent generation)—reveal these tradeoffs: gradient reversal achieves innocence but degrades general capabilities, representation manipulation preserves capabilities but generates incoherent text, and dataset filtering maintains coherence but fails to fully eliminate knowledge.
Through extensive experiments on WMDP biosecurity and cybersecurity benchmarks, as well as literary domain erasure, we measure the effectiveness of ELM’s principled approach to concept removal compared to previous methods. Our evaluation using both standard metrics and adversarial probing shows that ELM achieves robust concept erasure while maintaining model coherence and general capabilities.
@@ 2 Related work
Some techniques for controlling generative model outputs involve using an auxiliary generative model to specify the desired behavior over some training data, and then distilling these outputs into a target model by training the target model to mimic the demonstrated behavior. For example Askell et al. (2021) and Bai et al. (2022) prompt an unsafe model into behaving more safely, and then distil the prompted model. Similarly, Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. (Rosati et al., 2024) can also be classified under distillation methods as it mimics a Gaussian distribution when processing harmful tokens. Similarly, ELM matches the harmful logits to a modified output distribution, but unlike the distillation methods that usually adhere to the standard matching objective, we employ a multi-objective framework to address the seamlessness and specificity concerns with standard distillation methods. Prior works like Emulator and DeRa leverage probability ratios for behavioral modification - Emulator (Mitchell et al., 2023) scales fine-tuning changes between model sizes, while DeRa (Liu et al., 2024) controls alignment strength during inference. In contrast, ELM introduces a simpler, principled approach focused specifically on reducing the likelihood of knowledge concept generation.
@@
Gandikota et al. (2023) train a diffusion image model to mimic the outputs of an edited copy of the model whose generations have been guided to not produce images with certain attributes. Gandikota et al. (2024) erase concepts by modifying the key value mapping of cross attention layers in a low rank closed form update. Other works remove the knowledge of unwanted concepts from the model weights; proposing attention re-steering through fine-tuning (Zhang et al., 2023), fine-tuning the attention weights (Kumari et al., 2023) and continual learning (Heng & Soh, 2023). We take inspiration from (Gandikota et al., 2023) to reduce the likelihood of a concept being generated.
@@ 3.1 Transformer
Previous research (Meng et al., 2022) has localized much model knowledge within early to mid-layer blocks. As our own extensive experiments have similarly found good efficacy when erasing from earlier blocks of parameters, we restrict our parameter changes to MLP and MHA layers from those same blocks.
@@ 3.2 Transformer Next Token Prediction: A Classification Perspective
Language models are typically viewed through the lens of autoregressive sequence modeling, where the model predicts the next token given a sequence of previous tokens. However, these models can also be understood as powerful classifiers. To illustrate this connection, let’s examine both perspectives:
Traditional Perspective: The standard way to describe an autoregressive language model is through the factorization:
where the model predicts future tokens conditioned on previous tokens .
Classification Perspective: We can reformulate this using Bayes’ rule to obtain:
In this view, the model acts as a classifier where previous tokens serve as class labels for the text that follows. For instance, if we provide the context as ”The following text is about general biology”, the term represents the probability that the generated text belongs to the ”general biology” class.
This classification perspective enables us to manipulate the model’s knowledge by adjusting these classification probabilities. We can define a target distribution that modifies how strongly the model associates generated text with specific concepts:
where is a scaling factor that controls the strength of the modification. When , we increase the likelihood of generating text associated with the class ; when , we decrease it.
To implement this in practice with an autoregressive language model, we apply Bayes’ rule again:
In this form, both the numerator and denominator can be read out of the pretrained language model. Eq 6 is particularly powerful for concept unlearning because it allows us to use the language model itself as both the generator and classifier of content. For example, to unlearn bioweapons expertise, we can reduce the model’s probability of generating text that would naturally follow the context ”You are an expert in bioweapons.” This approach requires no external classifier - we leverage the model’s own understanding of concept relationships as encoded in its next-token predictions. Our perspective is inspired by the classifier-free guidance work (Ho & Salimans, 2022) in diffusion model research which has later been adopted to language models (Sanchez et al., 2023).
@@ 4 Method
We introduce Erasure of Language Memory (ELM), an approach that reformulates concept unlearning through introspective classification. While traditional unlearning methods have focused on sample removal through dataset retraining, gradient ascent, or representation disruption, ELM leverages the language model’s own ability to evaluate and modify its knowledge.
Specifically, ELM uses the language model itself as its own introspective classifier, as established in Section 3.2. Rather than training external classifiers or explicitly manipulating representations, we leverage the model’s next-token predictions to identify and reduce the likelihood of generating concept-specific content. This allows us to operate directly on the model’s probability distributions, providing a principled approach to unlearning.
The core of our method is a self-classification objective that reduces the likelihood of generating text that the model would classify as containing the target concept. Given a dataset containing text related to the target concept, we modify the model’s predicted probabilities to diverge from generating content related to the concept being erased and instead favor generating content related to a safer alternative :
where controls the strength of knowledge modification. is the concept to be erased, and is an alternative concept. The corresponding loss is:
In practice, we encounter two main challenges when implementing this objective. First, knowledge in language models is often entangled - modifying one concept can unintentionally affect related concepts. To address this, we preserve the model’s behavior on a set of related but safe concepts by matching its original distribution on retain data, , containing unrelated text:
Second, the self-classification objective alone might lead to incoherent text generation when prompted about erased concepts. We address this by applying our core objective (Equation 4) during inference to generate synthetic training examples, then using these to train the model in an autoregressive setting to maintain coherent generation:
The final training objective combines these terms with appropriate weights:
We implement this through low-rank adapters attached to early model layers, allowing precise modification of the model’s knowledge while maintaining its broader capabilities. This approach provides a principled method for concept unlearning that avoids the instability of gradient reversal methods and the incoherence of representation hacking approaches.
@@ 5.1 Experimental Setup
Our primary evaluation focuses on the Weapons of Mass Destruction Proxy (WMDP) dataset (Li et al., 2024), specifically utilizing the biosecurity (WMDP-bio) and cybersecurity (WMDP-cyber) multiple-choice questions (MCQs). To demonstrate ELM’s versatility, we also employ a modified version of the Harry Potter MCQ dataset (Lynch et al., 2024), expanded from binary to quaternary choices for consistency with other benchmarks. This diverse set of tasks allows us to assess ELM’s erasure effectiveness across different domains and knowledge types.
@@
We apply ELM to a range of state-of-the-art language models, including Zephyr-7B Beta (Tunstall et al., 2023b), Mistral-7B (Jiang et al., 2023), Llama3-8B, and Llama3-8B-instruct (Dubey et al., 2024) for the WMDP erasure tasks. For the Harry Potter knowledge erasure, we use the Llama-2-7B Chat model (Touvron et al., 2023) to maintain consistency with prior work from (Eldan & Russinovich, 2023). This selection of models enables us to evaluate ELM’s performance across various model architectures and training paradigms.
@@
From WMDP Bio forget corpus, we utilize 5,000 text samples, each with a maximum length of 700 characters. From Cyber forget corpus we use 1,000 texts of similar length. The Harry Potter erasure task employs 3,000 text samples extracted from the novel series, also limited to 700 characters each. To facilitate conditional erasure (Eq. 4), we prepend contexts such as “You are an expert in” followed by concept-specific keywords. Additionally, we incorporate text completion examples for consistency, following the approach used by (Qi et al., 2024).
@@
We assess our method, Erasure of Language Memory (ELM), along four key dimensions:
@@ 5.3 Ablation Study
We conduct ablation experiments to analyze the contribution of each loss component in ELM. Table 2 shows the impact on concept erasure (WMDP), general knowledge (MMLU), and generation quality (MT-Bench, Perplexity) when removing or modifying individual terms.
First, proves crucial for innocence. Removing significantly degrades erasure performance, with WMDP scores remaining close to the original model. While randomizing logits as a proxy for erasure (replacing in Eq. 10 with random vector) can achieve some erasure in cyber concepts, they lead to incoherent outputs.
@@
Ablating , leads to effective erasure but the model generates low-quality gibberish text when prompted for erased concept. Replacing it with random text from WikiText (Merity et al., 2016) slightly reduces fluency while maintaining erasure effectiveness. Qualitatively this produces awkward outputs that tend to be irrelevant to input prompts. This underscores the term’s role in maintaining seamless contextual relevance.
@@
The full ELM method achieves the best balance between concept erasure and general performance. We show a qualitative example with each of the settings below:
@@ 5.4 Specificity Analysis
To assess the specificity of our erasure method, we examine its impact on related MMLU classes. Figure 2 shows the performance of ELM and RMU on related safe concepts whose accuracies has to remain high (higher is better) when WMDP-bio and WMDP-cyber knowledge is erased. We find that both the methods reduces the accuracies slightly on closely-related safer concepts.